# Adversary Resilient Learned Bloom Filters

### v3

@@ 2.1 Classical Adversarial Model
As a running example, consider the set $S=\{x\}$ in a Classical Bloom Filter that uses $2$ hash functions, $h_{1},h_{2}$ and $4$ bits such that $h_{1}(x)=1$ and $h_{2}(x)=3$. A trivial deterministic setup algorithm, on input $S$, would then generate the representation $M^{S}_{r}=1010$.
Definition 2.2 (Query).
We are provided a set $S$ and a compressed representation of that set, $M_{r}^{S}$ (where $r$ is any random string). We define $Q_{s}$ to be a query algorithm such that $Q_{s}(M_{r}^{S},x)=1$ if $x\in S$, and $Q_{s}(M_{r}^{S},x)\in\{0,1\}$ if $x\notin S$. $Q_{s}$ may not be randomized and may not change $M_{r}^{S}$.
@@
Definition 2.4 (Classical Bloom Filter).
Let a Classical Bloom Filter be a data structure $B=(C_{r},Q)$ where $C_{r}$ obeys Definition 2.1 and $Q$ obeys either Definition 2.2 or Definition 2.3.
We define a special class of Classical Bloom Filters which were coined “steady” Classical Bloom Filters by Naor Yogev [NE19]. Steady Classical Bloom Filters do not change their internal representation $M^{S}_{r}$ after the setup algorithm $C_{r}$ has executed. In other words, only query algorithms of the type $Q_{s}$ are permitted in the steady setting and query algorithms of the type $Q_{u}$ are not permitted.
Definition 2.5 (Steady).
Let a steady $(n,\epsilon)$-Classical Bloom Filter be a Classical Bloom Filter $B_{s}=(C_{r},Q_{s})$ such that $Q_{s}$ obeys Definition 2.2 and $\forall x\in U$, it holds that
@@
Now we construct our first adversarial challenge for the Classical Bloom Filter in the steady setting. A probabilistic polynomial time (PPT) adversary $A$, as defined at the start of this section, is given a security parameter $1^{\lambda+n\log(u)}$ and is allowed to construct a set $S$. The set $S$ is then given to construction algorithm $C_{r}$ along with the security parameter to yield representation $M^{S}_{r}$. The adversary is allowed $t$ queries to the query algorithm $Q_{s}$ for which it is provided the results. After the $t$ queries, the adversary must output an element $x^{*}$. If $x^{*}$ is a false positive, and has not been queried before, the adversary wins the challenge. Otherwise, the adversary loses the challenge. We define this precisely in Challenge 1
Challenge 1 (Resilient).
@@
Now, we create an extension to Naor Yogev’s [NE19] model, introducing a stronger adversary that has access to the internal state of the Classical Bloom Filter. We construct our second adversarial challenge for the Classical Bloom Filter in the steady setting. This challenge is almost identical to the first challenge with the only difference being that the adversary is allowed access to the representation $M_{r}^{S}$. We define our second challenge more precisely in Challenge 2.
Challenge 2 (Reveal Resilient).
@@
    $A_{Q}$ is allowed access to $M_{r}^{S}$
@@
Analogous to Definition 2.6, we now define an adversarial reveal resilient Classical Bloom Filter based on the random variable $\Lambda^{2}_{A,t}(\lambda)$ (from Challenge 2).
Definition 2.7 (Reveal Resilient).
@@ 2.2 Learned Adversarial Model
Mitzenmacher [Mit18] was the first to create a mathematical model for the Learned Bloom Filter, in 2018. Our model is heavily based on Mitzenmacher’s model but with some additional definitions to suit our adversarial setting. Consider a set of elements $\mathcal{K}\subset S$ and a set of elements $\mathcal{U}$ such that $\forall u\in\mathcal{U},u\notin S$. We form a training dataset $\mathcal{D}=\{(x_{i},y_{i}=1)|x_{i}\in\mathcal{K}\}\cup\{(x_{i},y_{i}=0)|x_{i}\in\mathcal{U}\}$.
Definition 2.8 (Dataset Construction).
@@
Definition 2.9 (Learning Model).
@@
Definition 2.10 (Learned Construction).
@@
Definition 2.11 (Learned Query).
We define $\tilde{Q}_{s}$ to be a query algorithm similar to $Q_{s}$ (Definition  2.2) differing only in that $\tilde{Q}_{s}$ only takes a learned compressed representation $\tilde{M}_{r}^{S}$ of the set $S$ instead of any compressed representation $M_{r}^{S}$.
@@
Definition 2.12 (Learned Bloom Filter).
Let a Learned Bloom Filter be a data structure $\tilde{B}=(\tilde{C}_{r},\tilde{Q_{s}})$ where $\tilde{C}_{r}$ obeys Definition 2.10 and $\tilde{Q}_{s}$ obeys Definition 2.11. The query algorithm $\tilde{Q}_{s}$ for the Learned Bloom Filter is
@@
It is trivial to define a steady Learned Bloom Filter with completeness and soundness properties analogous to the ones outlined in Definition 2.5 for a Classical Bloom Filter. We now construct an adversarial model for the Learned Bloom Filter.
Challenge 3 (Learned Resilient).
@@
Note that the adversary may not choose threshold $\tau$. If the adversary is allowed to choose threshold $\tau$, then Challenge 3 is easily succeeded by choosing $\tau=0$. We are now ready to formally define security for the steady Learned Bloom Filter.
Definition 2.13 (Learned Resilient).
@@
Challenge 4 (Learned Reveal Resilient).
@@
    A is allowed access to $\tilde{M}_{r}^{S}$
@@
Definition 2.14 (Learned Reveal Resilient).
@@ 3 Secure Constructions for the Learned Bloom Filter
In this section, we propose a solution to the problem of securing the Learned Bloom Filter. We call our construction the Downtown Bodega Filter. Section 3.1 contains the background necessary to construct the Downtown Bodega Filter. We formally define the Downtown Bodega Filter in Section 3.2. The interested reader may also refer to Appendix A which discusses an alternate construction of a secure Learned Bloom Filter called the Uptown Bodega Filter.
@@ 3.1 Preliminaries
Setting Review: Consider a set of elements $\mathcal{K}\subset S$ and a set of elements $\mathcal{U}$ such that $\forall u\in\mathcal{U},u\notin S$. We form a dataset $\mathcal{D}=\{(x_{i},y_{i}=1)|x_{i}\in\mathcal{K}\}\cup\{(x_{i},y_{i}=0)|x_{i}\in\mathcal{U}\}$. We train a Bloom Filter Learning Model, $l_{r}^{\mathcal{D}}$, on the training dataset $\mathcal{D}$. Let $l_{r}^{\mathcal{D}}(x)$ be the probability estimate from the learning model (Definition 2.9) that $x$ is an element in $S$. A value $\tau$ may be chosen as a threshold. When $l_{r}^{\mathcal{D}}(x)\geq\tau$ then the Learned Bloom Filter considers $x$ to be an element of $S$. Otherwise, the Learned Bloom Filter passes $x$ on to the Backup Classical Bloom Filter.
We introduce a Learned Bloom Filter technique called “sandwiching” introduced by Mitzenmacher [Mit18]. We then discuss the use of a pseudo-random permutation on the Bloom Filter input set first proposed by Naor Yogev [NE19].
Definition 3.1 (Sandwiched Learned).
Let a Sandwiched Learned Bloom Filter, $SB_{r}=(\tilde{C}_{r},\tilde{Q_{s}})$ be a data structure where $\tilde{C}_{r}$ obeys Definition 2.10, and $\tilde{Q}_{s}$ obeys Definition 2.11.
@@ A Bloom Filter Learning Model $l_{r}^{\mathcal{D}}$ trained on $\mathcal{D}$
Figure 3 shows an example of a Sandwiched Learned Bloom Filter. In our running example, the setup algorithm for a Sandwiched Learned Bloom Filter creates the same Bloom Filter Learning Model, $l_{r}$, as well as the same compressed representation for the Backup Classical Bloom Filter $M_{r}^{S^{\prime}}$ (trained on the set $S^{\prime}=\{y\}$) that we discussed after Definition 2.10. In addition, it also creates a representation $M_{r}^{S}$ for the Initial Classical Bloom Filter trained on the complete set $S=\{x,y\}$. A trivial query algorithm $\tilde{Q}_{s}(\tilde{M}_{r}^{S},x)$ for the Sandwiched Learned Bloom Filter would then be $(M_{r}^{S}[h_{1}(x)]=1\land M_{r}^{S}[h_{2}(x)]=1)\land(l_{r}(x)>\tau\lor(M_{r}^{S^{\prime}}[h_{1}(x)]=1\land M_{r}^{S^{\prime}}[h_{2}(x)]=1)$.
Lemma 3.2.
@@
Proof 3.3.
The proof follows from the definition of a Sandwiched Learned Bloom Filter.
We use the standard definitions for pseudo-random permutations in this work. We provide a brief but self-contained treatment of pseudo-random permutations adapted from Chapter 3 of Katz Lindell [KL14] in Appendix B. Theorem $4.8$ of Naor Yogev [NE19] proves that for a classical steady $(n,\epsilon)$-Bloom Filter that uses $m$ bits of memory, if pseudo-random permutations exist, then there exists a negiligible function $\mathtt{negl}$ such that for security parameter $\lambda$ there exists a $(n,\epsilon+\mathtt{negl}(\lambda))$-adversarial resilient Classical Bloom Filter that uses $m^{{}^{\prime}}=m+\lambda$ bits of memory. This secure Classical Bloom Filter can be constructed by running the initialization algorithm on $S^{{}^{\prime}}={\{F_{k}(x):x\in S\}}$ instead of $S$ [NE19].
@@ 3.2 The Downtown Bodega Filter
Figure 4: A Downtown Bodega Filter. Both Positives and Negatives results for the Bloom Filter Learning Model are routed to Backup Bloom Filters secured with Pseudo-random Permutations
Definition 3.4 (Downtown).
@@
$DB_{r,k_{A},k_{B}}=(F_{k_{A}},F_{k_{B}},\tilde{C}_{r},\tilde{Q}_{s})$ where $\tilde{C}_{r}$ obeys Definition 2.10, $\tilde{Q}_{s}$ obeys Definition 2.11, and $F_{k_{A}},F_{k_{B}}$ are pseudo-random permutations. In other words, $(\tilde{C}_{r},\tilde{Q}_{s})$ form a Learned Bloom Filter.
The learned compressed representation of any set $S$ and training dataset $\mathcal{D}$ under an Downtown Bodega Filter consists of the following:
@@ A Bloom Filter Learning Model $l_{r}^{\mathcal{D}}$ trained on $\mathcal{D}$
Lemma 3.5 is just a rephrasing of Theorem $4.8$ of Naor Yogev [NE19] and follows directly from the theorem. We include a self-contained proof here for completeness. We also make it more evident that the proof holds not just for adversarial resilient Bloom Filters but also for adversarial reveal resilient Bloom Filters that provide guarantees under a strictly stronger adversarial model. It is important to note that, while the adversary has access to the internal representation of the Downtown Bodega Filter, the adversary does not have access to the secret keys $k_{A}$ and $k_{B}$. This is consistent with the formulation of Naor Yogev [NE19] who state at the end of Section 4 in their paper, with reference to Theorem $4.8$, “Notice that, in all the above constructions only the pseudo-random function (permutation) key must remain secret. That is, we get the same security even when the adversary gets the entire memory of the Bloom filter except for the PRF (PRP) key.”
Proof 3.6.
Let us consider Backup Classical Bloom Filter $A$ (a proof for Backup Bloom Filter $B$ can be constructed in exactly the same way). Our setup algorithm $\tilde{C}_{r}$ merely initializes Backup Classical Bloom Filter $A$ with $S^{\prime\prime}$. Our query algorithm on input $x$ queries for $x^{{}^{\prime}}=F_{k_{A}}(x)$. The only additional memory required is for storing $k_{A}$ which is $\lambda$ bits long.
The completeness follows from the completeness of the Classical Bloom Filter. The resilience of the construction follows from the following argument: consider an experiment where $F_{k_{A}}$ in Backup Classical Bloom Filter $A$ is replaced by a truly random oracle $\mathcal{R}(\cdot)$. Since $x$ has not been queried, we know that $R(x)$ is a truly random element that was not queried before, and we may think of it as chosen prior to the initialization of Backup Bloom Filter $A$. From the soundness of Backup Classical Bloom Filter $A$, we get that the probability of $x$ being a false positive is at most $\epsilon$.
@@
Proof 3.8.
We construct a Downtown Bodega Filter $DB_{r,k_{A},k_{B}}$ from a Sandwiched Learned Bloom Filter $SB_{r}$ as follows. We use the memory budget of the Initial Classical Bloom Filter to construct the Backup Classical Bloom Filter $A$. We use the memory budget of the Backup Classical Bloom Filter of the Sandwiched Learned Bloom Filter to construct Backup Classical Bloom Filter $B$. We do not modify the Bloom Filter Learning Model $l$ which remains trained on dataset $\mathcal{D}$ with threshold $\tau$. We choose keys $k_{A},k_{B}\in\{0,1\}^{\lambda}$ and use $2\lambda$ bits of extra memory to store them.
The completeness of $DB_{r,k_{A},k_{B}}$ follows from 1) the completeness of Backup Classical Bloom Filter $B$ and 2) the fact that any $x$ such that $l(x)<\tau$ is declared to be not in $S$ by $\tilde{Q}_{s}$ the query algorithm f the Downtown Bodega Filter if and only if the query algorithm of the Classical Bloom Filter with Backup Classical Bloom Filter $B$, $Q_{s}(M_{r}^{S^{\prime}},F_{k_{B}}(x))$, is also 00. This fact follows directly from the query algorithm $\tilde{Q}_{s}$ for the Downtown Bodega Filter stated in Definition 3.4.
To prove the resilience of the construction, we first show that the security of the Downtown Bodega Filter construction is reducible to the security of Backup Classical Bloom Filter $A$ and Backup Classical Bloom Filter $B$. Consider a false positive i.e an $x\notin S$ for which the Downtown Bodega Filter returns $1$. From the definition of the Downtown Bodega Filter, one of the following two cases must be true.
@@
Therefore, for any probabilistic polynomial time (PPT) adversary to induce a false positive in the overall Downtown Bodega Filter construction, they must either induce a false positive in Backup Classical Bloom Filter $A$ or Backup Classical Bloom Filter $B$. We have already proven in Lemma 3.5 that both Backup Classical Bloom Filter $A$ and Backup Classical Bloom Filter $B$ are $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient. It follows that the entire construction is $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient. This concludes our proof.

### v4

@@ 2.1 Classical Adversarial Model
As a running example, let $x$ be an element and consider the set $S=\{x\}$ in a Classical Bloom Filter that uses $2$ hash functions, $h_{1},h_{2}$ and $4$ bits such that $h_{1}(x)=1$ and $h_{2}(x)=3$. A trivial deterministic setup algorithm, on input $S$, would then generate the representation $M^{S}_{r}=1010$. The first and third bits in this representation are set according to the example.
Definition 6 (Query)
We are provided a set $S$ and a compressed representation of that set, $M_{r}^{S}$ (where $r$ is any random string). We define $Q_{s}$ to be a query algorithm such that $Q_{s}(M_{r}^{S},x)=1$ if $x\in S$, and $Q_{s}(M_{r}^{S},x)\in\{0,1\}$ if $x\notin S$. $Q_{s}$ must not be randomized and must not change $M_{r}^{S}$.
@@
Definition 8 (Classical Bloom Filter)
Let a Classical Bloom Filter be a data structure $B=(C_{r},Q)$ where $C_{r}$ obeys Definition 5 and $Q$ obeys either Definition 6 or Definition 7.
We define a special class of Classical Bloom Filters which were coined “steady” Classical Bloom Filters by Naor and Yogev [8]. Steady Classical Bloom Filters do not change their internal representation $M^{S}_{r}$ after the setup algorithm $C_{r}$ has been executed. In other words, only query algorithms of the type $Q_{s}$ are permitted in the steady setting, and query algorithms of the type $Q_{u}$ are not permitted.
Definition 9 (Steady)
Let a steady $(n,\epsilon)$-Classical Bloom Filter be a Classical Bloom Filter $B_{s}=(C_{r},Q_{s})$ such that $Q_{s}$ obeys Definition 6 and $\forall x\in U$, it holds that
@@
Now we construct our first adversarial challenge for the Classical Bloom Filter in the steady setting. A probabilistic polynomial time (PPT) adversary $A$, as defined at the start of this section, is given a security parameter $1^{\lambda+n\log(u)}$ and is allowed to construct a set $S$. The set $S$ is then given to construction algorithm $C_{r}$ along with the security parameter to yield representation $M^{S}_{r}$. The adversary is allowed $t$ queries to the query algorithm $Q_{s}$ for which it is provided the results. After the $t$ queries, the adversary must output an element $x^{*}$. If $x^{*}$ is a false positive and has not been queried before, the adversary wins the challenge. Otherwise, the adversary loses the challenge. We define this precisely in Challenge 2.1
Challenge 2.1 (Resilient)
@@
Now, we create an extension to Naor and Yogev ’s [8] model, introducing a stronger adversary that has access to the internal state of the Classical Bloom Filter. We construct our second adversarial challenge for the Classical Bloom Filter in the steady setting. This challenge is almost identical to the first challenge with the only difference being that the adversary is allowed access to the representation $M_{r}^{S}$. We define our second challenge more precisely in Challenge 2.2.
Challenge 2.2 (Reveal Resilient)
@@
    $x^{*}\leftarrow A_{Q}^{Q_{s}(M^{S}_{r},\cdot)}(1^{\lambda+n\log(u)},S,M_{r}^{S})$. $A_{Q}$ performs at most $t$ queries $x_{1},\ldots,x_{t}$ to $Q_{s}(M^{S}_{r},\cdot)$.
@@
Analogous to Definition 10, we now define an adversarial reveal resilient Classical Bloom Filter based on the random variable $\Lambda^{2}_{A,t}(\lambda)$ (from Challenge 2.2).
Definition 11 (Reveal Resilient)
@@ 2.2 Learned Adversarial Model
Mitzenmacher [7] was the first to create a mathematical model for the Learned Bloom Filter, in 2018. Our model is heavily based on Mitzenmacher’s model but with some additional definitions to suit our adversarial setting. Consider a set of elements $\mathcal{K}\subset S$ and a set of elements $\mathcal{U}$ such that $\forall u\in\mathcal{U},u\notin S$. We form a training dataset $\mathcal{D}=\{(x_{i},y_{i}=1)|x_{i}\in\mathcal{K}\}\cup\{(x_{i},y_{i}=0)|x_{i}\in\mathcal{U}\}$.
Definition 12 (Dataset Construction)
@@
Definition 13 (Learning Model)
@@
Definition 14 (Learned Construction)
@@
Definition 15 (Learned Query)
We define $\tilde{Q}_{s}$ to be a query algorithm similar to $Q_{s}$ (Definition  6) differing only in that $\tilde{Q}_{s}$ only takes a learned compressed representation $\tilde{M}_{r}^{S}$ of the set $S$ instead of any compressed representation $M_{r}^{S}$.
@@
Definition 16 (Learned Bloom Filter)
Let a Learned Bloom Filter be a data structure $\tilde{B}=(\tilde{C}_{r},\tilde{Q_{s}})$ where $\tilde{C}_{r}$ obeys Definition 14 and $\tilde{Q}_{s}$ obeys Definition 15. The query algorithm $\tilde{Q}_{s}$ for the Learned Bloom Filter is
@@
It is trivial to define a steady Learned Bloom Filter with completeness and soundness properties analogous to the ones outlined in Definition 9 for a Classical Bloom Filter. We now construct an adversarial model for the Learned Bloom Filter.
Challenge 2.3 (Learned Resilient)
@@
Note that the adversary must not choose threshold $\tau$. If the adversary is allowed to choose threshold $\tau$, then Challenge 2.3 is easily succeeded by choosing $\tau=0$. We are now ready to formally define security for the steady Learned Bloom Filter.
Definition 17 (Learned Resilient)
@@
Challenge 2.4 (Learned Reveal Resilient)
@@
    $x^{*}\leftarrow A_{Q}^{\tilde{Q}_{s}(\tilde{M}^{S}_{r},\cdot)}(1^{\lambda+n\log(u)},S,\tilde{M}_{r}^{S})$. $A_{Q}$ performs at most $t$ queries $x_{1},\ldots,x_{t}$ to $\tilde{Q}_{s}(\tilde{M}^{S}_{r},\cdot)$.
@@
Definition 18 (Learned Reveal Resilient)
@@ 3 Secure Constructions for the Learned Bloom Filter
In this section, we propose a solution to the problem of securing the Learned Bloom Filter. We call our construction the Downtown Bodega Filter. Section 3.1 contains the background necessary to construct the Downtown Bodega Filter. We formally define the Downtown Bodega Filter in Section 3.2.
@@ 3.1 Preliminaries
Setting Review: Consider a set of elements $\mathcal{K}\subset S$ and a set of elements $\mathcal{U}$ such that $\forall u\in\mathcal{U},u\notin S$. We form a dataset $\mathcal{D}=\{(x_{i},y_{i}=1)|x_{i}\in\mathcal{K}\}\cup\{(x_{i},y_{i}=0)|x_{i}\in\mathcal{U}\}$. We train a Bloom Filter Learning Model, $l_{r}^{\mathcal{D}}$, on the training dataset $\mathcal{D}$. Let $l_{r}^{\mathcal{D}}(x)$ be the probability estimate from the learning model (Definition 13) that $x$ is an element in $S$. A value $\tau$ may be chosen as a threshold. When $l_{r}^{\mathcal{D}}(x)\geq\tau$ then the Learned Bloom Filter considers $x$ to be an element of $S$. Otherwise, the Learned Bloom Filter passes $x$ onto the Backup Classical Bloom Filter.
We introduce a Learned Bloom Filter technique called “sandwiching” introduced by Mitzenmacher [7]. We then discuss the use of a pseudo-random permutation on the Bloom Filter input set first proposed by Naor and Yogev [8].
Definition 19 (Sandwiched Learned)
Let a Sandwiched Learned Bloom Filter, $SB_{r}=(\tilde{C}_{r},\tilde{Q_{s}})$ be a data structure where $\tilde{C}_{r}$ obeys Definition 14, and $\tilde{Q}_{s}$ obeys Definition 15.
@@ A Bloom Filter Learning Model $l_{r}^{\mathcal{D}}$ trained on $\mathcal{D}$
Figure 3 shows an example of a Sandwiched Learned Bloom Filter. In our running example, the setup algorithm for a Sandwiched Learned Bloom Filter creates the same Bloom Filter Learning Model, $l_{r}$, as well as the same compressed representation for the Backup Classical Bloom Filter $M_{r}^{S^{\prime}}$ (trained on the set $S^{\prime}=\{y\}$) that we discussed after Definition 14. In addition, it also creates a representation $M_{r}^{S}$ for the Initial Classical Bloom Filter trained on the complete set $S=\{x,y\}$. A trivial query algorithm $\tilde{Q}_{s}(\tilde{M}_{r}^{S},x)$ for the Sandwiched Learned Bloom Filter would then be $(M_{r}^{S}[h_{1}(x)]=1\land M_{r}^{S}[h_{2}(x)]=1)\land(l_{r}(x)>\tau\lor(M_{r}^{S^{\prime}}[h_{1}(x)]=1\land M_{r}^{S^{\prime}}[h_{2}(x)]=1)$.
Lemma 1
@@
Proof
The proof follows from the definition of a Sandwiched Learned Bloom Filter. $x$ is a false positive in one of two cases:
Case 1: The Learning Model returns positive on $x$. In this case, $x$ only reaches the Learning Model if it was not marked negative by the Initial Classical Bloom Filter. Therefore it was a false positive in the Initial Classical Bloom Filter.
Case 2: The Backup Classical Bloom Filter returns positive on $x$. Using similar reasoning as Case 1, we can show that this only occurs if $x$ was a false positive in the Initial Classical Bloom Filter.
We use the standard definitions for pseudo-random permutations in this work. We provide a brief but self-contained treatment of pseudo-random permutations adapted from Chapter 3 of Katz Lindell [14] in Appendix 0.A. Theorem $4.8$ of Naor and Yogev [8] proves that for a classical steady $(n,\epsilon)$-Bloom Filter that uses $m$ bits of memory, if pseudo-random permutations exist, then there exists a negligible function $\mathtt{negl}$ such that for security parameter $\lambda$ there exists a $(n,\epsilon+\mathtt{negl}(\lambda))$-adversarial resilient Classical Bloom Filter that uses $m^{{}^{\prime}}=m+\lambda$ bits of memory. This secure Classical Bloom Filter can be constructed by running the initialization algorithm on $S^{{}^{\prime}}={\{F_{k}(x):x\in S\}}$ instead of $S$ [8].
@@ 3.2 The Downtown Bodega Filter
Figure 4: A Downtown Bodega Filter. Both Positive and Negative results for the Bloom Filter Learning Model are routed to Backup Bloom Filters secured with Pseudo-random Permutations
Definition 20 (Downtown)
@@
$DB_{r,k_{A},k_{B}}=(F_{k_{A}},F_{k_{B}},\tilde{C}_{r},\tilde{Q}_{s})$ where $\tilde{C}_{r}$ obeys Definition 14, $\tilde{Q}_{s}$ obeys Definition 15, and $F_{k_{A}},F_{k_{B}}$ are pseudo-random permutations. In other words, $(\tilde{C}_{r},\tilde{Q}_{s})$ form a Learned Bloom Filter.
The learned compressed representation of any set $S$ and training dataset $\mathcal{D}$ under a Downtown Bodega Filter consists of the following:
@@ A Bloom Filter Learning Model $l_{r}^{\mathcal{D}}$ trained on $\mathcal{D}$
Lemma 2 is just a rephrasing of Theorem $4.8$ of Naor and Yogev [8] and follows directly from the theorem. We include a self-contained proof here for completeness. We also make it more evident that the proof holds not just for adversarial resilient Bloom Filters but also for adversarial reveal resilient Bloom Filters that provide guarantees under a strictly stronger adversarial model. It is important to note that, while the adversary has access to the internal representation of the Downtown Bodega Filter, the adversary does not have access to the secret keys $k_{A}$ and $k_{B}$. This is consistent with the formulation of Naor and Yogev [8] who state at the end of Section 4 in their paper, concerning Theorem $4.8$, “Notice that, in all the above constructions only the pseudo-random function (permutation) key must remain secret. That is, we get the same security even when the adversary gets the entire memory of the Bloom filter except for the PRF (PRP) key.”
Proof
Let us consider Backup Classical Bloom Filter $A$ (a proof for Backup Bloom Filter $B$ can be constructed in the same way). Our setup algorithm $\tilde{C}_{r}$ merely initializes Backup Classical Bloom Filter $A$ with $S^{\prime\prime}$. Our query algorithm on input $x$ queries for $x^{{}^{\prime}}=F_{k_{A}}(x)$. The only additional memory required is for storing $k_{A}$ which is $\lambda$ bits long.
The completeness follows from the completeness of the Classical Bloom Filter. The resilience of the construction follows from the following argument: consider an experiment where $F_{k_{A}}$ in Backup Classical Bloom Filter $A$ is replaced by a truly random oracle $\mathcal{R}(\cdot)$. Since $x$ has not been queried, we know that $R(x)$ is a truly random element that was not queried before, and we may think of it as chosen before the initialization of Backup Bloom Filter $A$. From the soundness of Backup Classical Bloom Filter $A$, we get that the probability of $x$ being a false positive is at most $\epsilon$.
@@
Proof
We construct a Downtown Bodega Filter $DB_{r,k_{A},k_{B}}$ from a Sandwiched Learned Bloom Filter $SB_{r}$ as follows. We use the memory budget of the Initial Classical Bloom Filter to construct the Backup Classical Bloom Filter $A$. We use the memory budget of the Backup Classical Bloom Filter of the Sandwiched Learned Bloom Filter to construct the Backup Classical Bloom Filter $B$. We do not modify the Bloom Filter Learning Model $l$ which remains trained on dataset $\mathcal{D}$ with threshold $\tau$. We choose keys $k_{A},k_{B}\in\{0,1\}^{\lambda}$ and use $2\lambda$ bits of extra memory to store them.
The completeness of $DB_{r,k_{A},k_{B}}$ follows from 1) the completeness of Backup Classical Bloom Filter $B$ and 2) the fact that any $x$ such that $l(x)<\tau$ is declared to be not in $S$ by $\tilde{Q}_{s}$ the query algorithm f the Downtown Bodega Filter if and only if the query algorithm of the Classical Bloom Filter with Backup Classical Bloom Filter $B$, $Q_{s}(M_{r}^{S^{\prime}},F_{k_{B}}(x))$, is also 00. This fact follows directly from the query algorithm $\tilde{Q}_{s}$ for the Downtown Bodega Filter stated in Definition 20.
To prove the resilience of the construction, we first show that the security of the Downtown Bodega Filter construction is reducible to the security of Backup Classical Bloom Filter $A$ and Backup Classical Bloom Filter $B$. Consider a false positive i.e. an $x\notin S$ for which the Downtown Bodega Filter returns $1$. From the definition of the Downtown Bodega Filter, one of the following two cases must be true.
@@
Therefore, for any probabilistic polynomial time (PPT) adversary to induce a false positive in the overall Downtown Bodega Filter construction, they must either induce a false positive in Backup Classical Bloom Filter $A$ or Backup Classical Bloom Filter $B$. We have already proven in Lemma 2 that both Backup Classical Bloom Filter $A$ and Backup Classical Bloom Filter $B$ are $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient. It follows that the entire construction is $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient. This concludes our proof.